# Towards Seamless Adaptation of Pre-trained Models for Visual Place Recognition

### 22 Feb 2024 v1

@@ 1 Introduction
Recently, foundation models (Radford et al., 2021; Yuan et al., 2021; Oquab et al., 2023) achieved remarkable performance on many computer vision tasks given their ability to produce well-generalized representations. However, the image representation produced by the pre-trained model is susceptible to useless (even harmful) dynamic objects (e.g. pedestrians and vehicles), and tends to ignore some static discriminative backgrounds (e.g. buildings and vegetation), as shown in Fig. 1 (b). A robust VPR model should focus on the static discriminative landmarks (Chen et al., 2017b) rather than the dynamic foreground. This results in a gap between pre-training and VPR tasks. Meanwhile, full fine-tuning the foundation model on downstream datasets might forget previously learned knowledge and damage the excellent transferability, i.e., catastrophic forgetting. An effective method to address this issue is parameter-efficient transfer learning (Houlsby et al., 2019; Lester et al., 2021), which has not been studied in the VPR area. Besides, most foundation models do not directly produce local features, which is typically required in the re-ranking of two-stage VPR methods.
@@
Figure 1: Attention map visualizations of the pre-trained foundation model (DINOv2) and our model. The pre-trained model pays attention to some regions (e.g. dynamic pedestrians) that are useless to identify places. Our method focuses on discriminative regions (buildings and trees).
@@ 2 Related Work
Visual Place Recognition: The traditional VPR approaches perform nearest neighbor search using global features to find the most similar place. The global features are commonly produced using aggregation algorithms, such as Bag of Words (Angeli et al., 2008) and VLAD (Jégou et al., 2010), to process the hand-crafted features like SURF (Bay et al., 2008). As the advancement of deep learning techniques, many works (Sünderhauf et al., 2015; Arandjelovic et al., 2016; Jin Kim et al., 2017; Chen et al., 2017a; b; Naseer et al., 2017; Garg et al., 2017; 2018; Xin et al., 2019; Yin et al., 2019; Lu et al., 2021; Leyva-Vallina et al., 2021; 2023; Ali-Bey et al., 2023) employed a variety of deep features for the VPR task. Some works integrated the aggregation methods into neural networks (Arandjelovic et al., 2016; Yu et al., 2019; Peng et al., 2021), and improved training strategies (Ge et al., 2020; Berton et al., 2022a; Ali-bey et al., 2022), to achieve better performance. Nevertheless, most one-stage (i.e. global retrieval) approaches are prone to perceptual aliasing due to the use of aggregated features while neglecting spatial information. One recent work (Keetha et al., 2023) also used pre-trained foundation models for VPR and got promising results. However, this work did not perform any fine-tuning, making it difficult to fully unleash the capability of these models for VPR.
Recently, the two-stage (i.e. hierarchical) VPR methods (Hausler & Milford, 2020; Garg & Milford, 2021; Hausler et al., 2021; Berton et al., 2021; Shen et al., 2022; Wang et al., 2022a; Lu et al., 2023; Shen et al., 2023; Zhu et al., 2023) have become popular. These approaches typically retrieved top-k candidate images over the whole database using compact global feature representation (such as NetVLAD (Arandjelovic et al., 2016) or Generalized Mean (GeM) pooling (Radenović et al., 2018)), then re-ranked candidates by performing local matching between the query image and each candidate using local descriptors. However, most of these methods required geometric consistency verification after local matching (Cao et al., 2020; Hausler et al., 2021; Wang et al., 2022a) or taking into account spatial constraints during matching (Berton et al., 2021; Shen et al., 2022; Zhu et al., 2023), which greatly increases the runtime burden. The recent visual foundation model (Oquab et al., 2023) has shown an excellent ability to match similar semantic patch-level features across domains. In this work, we attempt to match local features produced by the foundation model for re-ranking without spatial verification, thereby significantly speeding up the retrieval process in VPR.
@@ 3.1 Preliminary
Given an input image, ViT first slices it into $N$ patches and linearly projects them to D-dim patch embeddings $x_{p}\in\mathcal{R}^{N\times D}$, then prepends a learnable [class] token to $x_{p}$ as $x_{0}=[x_{class};x_{p}]\in\mathcal{R}^{(N+1)\times D}$. After adding positional embeddings to preserve the positional information, $x_{0}$ is fed into a series of transformer blocks to produce the feature representation. A standard transformer block mainly includes multi-head attention (MHA), multi-layer perceptron ( MLP), and LayerNormalization (LN) layers, as shown in Fig. 2 (a). For the input token sequence, its change process passing through a transformer block is: The MHA is first applied to compute attentional features, then MLP is utilized to realize the feature nonlinearization and dimension transformation. It is formalized as:
@@
$x_{l}=\text{MLP}(\text{LN}(x^{\prime}_{l}))+x^{\prime}_{l}$ (2)
@@
For the feature map output by CNN models, a common practice of the two-stage VPR method is to use NetVLAD or GeM to aggregate it into a global feature for candidate retrieval and also treat it as dense local (patch-level) features for re-ranking. For the ViT model, the output consists of one class token and $N$ patch tokens, where the class token can be directly used as the global feature to represent places. Meanwhile, $N$ patch tokens can also be reshaped as a feature map (similar to CNN) to restore spatial position. In this work, instead of using the class token as the global feature, we use GeM to pool the feature map into the global feature. The reason is explained in Appendix B.
@@ 3.2 Global Adaptation
Although pre-trained foundation models are capable of powerful feature representation, direct use of them in VPR cannot fully unleash their capability due to the gap between pre-training and VPR tasks. To address it, we introduce the global adaptation to adapt the pre-trained model so that the feature representation can focus on the static discriminative regions that are beneficial to VPR.
@@
The output of the last transformer block is fed into an LN layer as the final output of the entire global adapted ViT backbone. We discard the class token and reshape patch tokens as the produced feature map $fm$. The L2-normalized GeM global feature used to retrieve candidates can be denoted as
@@
The global adapted foundation model can produce feature representations that focus on discriminative landmarks and ignore dynamic interference. This bridges the gap between pre-training and VPR tasks, and greatly boosts the performance of visual foundation models in the VPR task.
@@ 3.3 Local Adaptation
Two-stage VPR methods typically match dense local features to re-rank candidate places for boosting performance. In the above introduction, we have obtained the feature map output by global adapted ViT backbone, which can also be regarded as coarse-grained patch-level features. However, in order to achieve appreciable performance in local matching, more fine-grained dense local features are required. To achieve this, we propose the local adaptation, which is achieved by an up-sampling module after the ViT backbone (as shown in Fig .3). To be specific, this module consists of two up-convolutional (up-conv) layers and a ReLU layer in the middle. The height and width of the feature map will approximately double after passing through each up-conv layer, while the channel dimension will be reduced. Finally, for the input image with sizes of 224$\times$224, this module adjusts the 16$\times$16$\times$1024 feature map output by ViT-L/14 to 61$\times$61$\times$128, and performs L2 normalization in the channel dimension (intraL2) to yield dense local features, i.e., a dense 61$\times$61 grid of 128-dim local features $\bm{f}^{l}$. Formally, it can be represented as
@@ 3.4 Local Matching for Re-ranking
Our two-stage place retrieval pipeline is shown in Fig 3. After obtaining the global features and local features, we first compute L2 distance to perform the similarity search in the global feature space over the database to get the top-k most similar candidate images. For the local features matching between the query image $q$ and a candidate image $c$, we search for mutual nearest neighbor matches by cross-matching. Since local features are L2 normalized, the inner product equivalent to cosine similarity is used to measure local feature similarity. That is
$\bm{s}_{qc}(i,j)=\bm{f}^{l}_{q}(i)^{T}\bm{f}^{l}_{c}(j)\quad i,j\in N,$ (7)
@@
That is, $u$-th feature in image $q$ and $v$-th feature in image $c$ are the best matches for each other. The obtained nearest neighbor matches set in previous VPR work (Hausler et al., 2021; Wang et al., 2022a) exists a large number of false matches. So they apply geometric verification (e.g. RANSAC (Fischler & Bolles, 1981)) to remove outliers (i.e. false matches) and use the number of inliers as the image similarity score, which is time-consuming. Due to the powerful representation ability of the foundation model and the superiority of the ViT in capturing long-distance feature dependencies, the number of false matches in this work is actually not enough to affect the performance of re-ranking. So we directly use the number of matches (i.e., $|\mathcal{M}|$) as the image similarity score for re-ranking.
@@ 4.1 Datasets and Performance Evaluation
Several VPR benchmark datasets mainly including Tokyo24/7, MSLS, and Pitts30k are used in our experiments. Table 1 summarizes their main information. Tokyo24/7 (Torii et al., 2015) includes about 76k database images and 315 query images captured from urban scenes with drastic illumination changes. Mapillary Street-Level Sequences (MSLS) (Warburg et al., 2020) consists of more than 1.6 million images collected in urban, suburban and natural scenes over 7 years. We assess models on both MSLS-val and MSLS-challenge (an online test set without released labels) sets. Pittsburgh (Pitts30k) (Torii et al., 2013) contains 30k reference images and 24k query images in the train, val and test sets, and exists severe viewpoint changes. More details are in Appendix H.
@@ 4.2 Implementation Details
We use the DINOv2 based on ViT-L/14 as the foundation model and conduct all experiments on an NVIDIA GeForce RTX 3090 GPU using PyTorch. Fed a 224×224 image, the model produces a 1024-dim global feature and a dense grid of 128-dim local features. The bottleneck ratio of the adapters in ViT blocks is 0.5 and the scaling factor $s$ in Eq. 4 is set to 0.2. We use 3×3 up-conv with stride=2 and padding=1 in the local adaptation module. The output channels of the first and second up-conv layers are 256 and 128, respectively. Following other two-stage methods, we re-rank the top-100 candidates to yield final results. We train our models using the Adam optimizer with the learning rate set as 0.00001 and batch size set as 4. When the R@5 on the validation set does not have improvement within 3 epochs, the training is terminated. For MSLS, we set an epoch as passing 30k queries, whereas Pitts30k is passing 5k queries. In model training, we define the potential positive images as the reference images that are within 10 meters from the query image, while the definite negative images are those further than 25 meters. Two hard negative images from 1000 randomly chosen definite negatives are used in the triplet loss. We empirically set the margin $m=0.1$ in Eq. 9, the weight $\lambda=1$ in Eq. 11. For the experiments in Section 4.3, we fine-tune our models on MSLS-train to test on MSLS-val/challenge, and further fine-tune them on Pitts30k-train to test on Pitts30k-test and Tokyo24/7 (as in $R^{2}$ormer (Zhu et al., 2023)). For the experiment in other sections, the models tested on Pitts30k-test and Tokyo24/7 are directly fine-tuned on Pitts30k-train.
@@ 4.3 Comparisons with State-of-the-Art Methods
In this section, we compare the proposed SelaVPR method with several state-of-the-art (SOTA) VPR methods, including four one-stage methods using global feature retrieval: NetVLAD (Arandjelovic et al., 2016), SFRS (Ge et al., 2020), CosPlace (Berton et al., 2022a) and MixVPR (Ali-Bey et al., 2023), as well as five two-stage methods with re-ranking: SP-SuperGlue (DeTone et al., 2018; Sarlin et al., 2020), Patch-NetVLAD (Hausler et al., 2021), TransVPR (Wang et al., 2022a), StructVPR (Shen et al., 2023) and $R^{2}$ormer (Zhu et al., 2023). The details of these methods are in Appendix I. Note that CosPlace and MixVPR are trained on individually constructed large-scale datasets. TransVPR and $R^{2}$ormer are transformer-based methods. $R^{2}$ormer ranked first in the leaderboard of the MSLS place recognition challenge before our method was proposed. We also show the global retrieval result without re-ranking using the proposed SelaVPR, and denote it as SelaVPR(global). The quantitative results are shown in Table 2. Since the code for StructVPR and $R^{2}$ormer has not yet been released, we use the results reported in their paper. The proposed method achieves the best R@1/R@5/R@10 on all datasets.
When using only global features for direct retrieval, SelaVPR(global) significantly outperforms other one-stage methods on all datasets for R@5 and R@10, including SFRS using complex training strategies and CosPlace/MixVPR trained on purpose-built large-scale datasets. SelaVPR(global) also outperforms all other two-stage methods on Tokyo24/7, MSLS-val, and MSLS-challenge for R@5 and R@10. This fully demonstrates that adapting the foundation model can provide powerful feature representation, which is a novel way to achieve SOTA one-stage VPR. Although SelaVPR(global) does not achieve very good R@1 on Tokyo24/7 where the lighting changes drastically, the complete SelaVPR method outperforms other methods by a large margin after local feature re-ranking. This illustrates the necessity of using local matching (local adaptation) for VPR in extreme environments. The complete SelaVPR method significantly outperforms SOTA methods on Tokyo24/7 and Pitts30k with absolute R@1 improvement of 5.4% and 1.3% respectively. Meanwhile, it also ranks 1st on the leaderboard of MSLS place recognition challenge (see Appendix J). Comparisons to SOTA methods on more datasets are in Appendix D (SelaVPR achieves near-perfect results on St. Lucia). Fig. 4 qualitatively demonstrates that our approach is highly robust in challenging scenes. Benefiting from the visual foundation model and sensible adaptation, SelaVPR achieves absolute performance advantages on a variety of datasets without complex training strategies or purpose-built large-scale training datasets.
@@
Efficiency is another metric for evaluating the VPR methods. In Table 3, we compare the runtime (feature extraction time and matching/retrieval time) of our method with other two-stage methods on Pitts30k-test. Patch-NetVLAD-p and TransVPR are representative two-stage methods using RANSAC for spatial verification in re-ranking. SP-SuperGlue uses neural networks to match local features. TransVPR is fast at extracting features, while SelaVPR is slower (but faster than other methods) due to the use of the ViT/L backbone. However, for feature matching (i.e. retrieval), since our method does not require time-consuming spatial verification, its runtime is less than 3% of TransVPR and only about 1% of SP-SuperGlue and Patch-NetVLAD-p. The total runtime of our method is less than 4% of TransVPR. Although Patch-NetVLAD-s uses a Rapid Spatial Scoring method for fast verification, it is still significantly slower than ours. Fig. 5 simultaneously shows the total runtime and R@1 on Pitts30k. Due to the absolute advantages in both performance and efficiency, our SelaVPR method is able to pave the way for real-world large-scale VPR applications.

### 18 Mar 2024 v2

@@ 1 Introduction
Recently, foundation models (Radford et al., 2021; Yuan et al., 2021; Oquab et al., 2023) achieved remarkable performance on many computer vision tasks given their ability to produce well-generalized representations. However, the image representation produced by the pre-trained model is susceptible to useless (even harmful) dynamic objects (e.g. pedestrians and vehicles), and tends to ignore some static discriminative backgrounds (e.g. buildings and vegetation), as shown in Fig. 1 (b). A robust VPR model should focus on the static discriminative landmarks (Chen et al., 2017b) rather than the dynamic foreground. This results in a gap between pre-training and VPR tasks. Meanwhile, full fine-tuning the foundation model on downstream datasets might forget previously learned knowledge and damage the excellent transferability, i.e., catastrophic forgetting. An effective method to address this issue is parameter-efficient transfer learning (Houlsby et al., 2019; Lester et al., 2021), which has not been studied in the VPR area. Besides, most foundation models do not directly produce (dense) local features, which is typically required in the re-ranking of two-stage VPR methods.
@@
Figure 1: Attention map visualizations of the pre-trained foundation model (DINOv2) and our model. The pre-trained model pays attention to some regions (e.g. dynamic riders) that are useless to identify places. Our method focuses on discriminative regions (buildings and trees).
@@ 2 Related Work
Visual Place Recognition: The traditional VPR approaches perform nearest neighbor search using global features to find the most similar place. The global features are commonly produced using aggregation algorithms, such as Bag of Words (Angeli et al., 2008) and VLAD (Jégou et al., 2010), to process the hand-crafted features like SURF (Bay et al., 2008). With the advancement of deep learning techniques, many works (Sünderhauf et al., 2015; Jin Kim et al., 2017; Chen et al., 2017a; b; Naseer et al., 2017; Garg et al., 2017; 2018; Xin et al., 2019; Yin et al., 2019; Lu et al., 2021; Leyva-Vallina et al., 2021; 2023; Ali-Bey et al., 2023) have employed a variety of deep features for the VPR task. Some works integrated the aggregation methods into neural networks (Arandjelovic et al., 2016; Yu et al., 2019; Peng et al., 2021), and improved training strategies (Ge et al., 2020; Berton et al., 2022a; Ali-bey et al., 2022), to achieve better performance. Nevertheless, most one-stage (i.e. global retrieval) VPR approaches are prone to perceptual aliasing due to the use of aggregated features while neglecting spatial information. One recent work (Keetha et al., 2023) also used pre-trained foundation models for the VPR task. However, this work did not perform any fine-tuning, making it difficult to fully unleash the capability of these models for VPR.
Recently, the two-stage (i.e. hierarchical) VPR methods (Hausler & Milford, 2020; Garg & Milford, 2021; Hausler et al., 2021; Berton et al., 2021; Shen et al., 2022; Wang et al., 2022a; Lu et al., 2023; Shen et al., 2023; Zhu et al., 2023) have become popular. These approaches typically retrieved top-k candidate images over the whole database using compact global feature representation, such as NetVLAD (Arandjelovic et al., 2016) or Generalized Mean (GeM) pooling (Radenović et al., 2018), then re-ranked candidates by performing local matching between the query image and each candidate using local descriptors. However, most of these methods required geometric consistency verification after local matching (Cao et al., 2020; Hausler et al., 2021; Wang et al., 2022a) or taking into account spatial constraints during matching (Berton et al., 2021; Lu et al., 2023; Zhu et al., 2023), which greatly increases the runtime burden. The recent visual foundation model (Oquab et al., 2023) has shown an excellent ability to match similar semantic patch-level features across domains. In this work, we attempt to match local features produced by the foundation model for re-ranking without spatial verification, thereby significantly speeding up the retrieval process in VPR.
@@ 3.1 Preliminary
Given an input image, ViT first slices it into $N$ patches and linearly projects them to $D$-dim patch embeddings $x_{p}\in\mathcal{R}^{N\times D}$, then prepends a learnable [class] token to $x_{p}$ as $x_{0}=[x_{class};x_{p}]\in\mathcal{R}^{(N+1)\times D}$. After adding positional embeddings to preserve the positional information, $x_{0}$ is fed into a series of transformer blocks to produce the feature representation. A standard transformer block mainly includes multi-head attention (MHA), multi-layer perceptron ( MLP), and LayerNormalization (LN) layers, as shown in Fig. 2 (a). For the input token sequence, its change process passing through a transformer block is: The MHA is first applied to compute attentional features, then MLP is utilized to realize the feature nonlinearization and dimension transformation. It is formulated as:
@@
$x_{l}=\text{MLP}(\text{LN}(x^{\prime}_{l}))+x^{\prime}_{l},$ (2)
@@
For the feature map output by CNN models, a common practice of the two-stage VPR method is to use NetVLAD or GeM to aggregate it into a global feature for candidate retrieval and also treat it as dense local (patch) features for re-ranking. For the ViT model, the output consists of one class token and $N$ patch tokens, where the class token can be directly used as the global feature to represent places. Meanwhile, $N$ patch tokens can also be reshaped as a feature map (similar to CNN) to restore spatial position. In this work, instead of using the class token as the global feature, we use GeM to pool the feature map into the global feature. The reason is explained in Appendix C.
@@ 3.2 Global Adaptation
Although pre-trained foundation models are capable of powerful feature representation, direct use of them in VPR cannot fully unleash their capability due to the gap between the pre-training and VPR tasks. To address it, we introduce the global adaptation to adapt the pre-trained model so that the feature representation can focus on the static discriminative regions that are beneficial to VPR.
@@
The output of the last transformer block is fed into an LN layer as the final output of the entire global adapted ViT backbone. We discard the class token and reshape patch tokens as the produced feature map $fm$. The L2-normalized GeM global feature used to retrieve candidates can be written as
@@
The global adapted foundation model can produce feature representations that focus on discriminative landmarks and ignore dynamic interference. This bridges the gap between the model pre-training and VPR tasks, and greatly boosts the performance of foundation models in the VPR task.
@@ 3.3 Local Adaptation
Two-stage VPR methods typically match dense local features to re-rank candidate places for boosting performance. In the above introduction, we have obtained the feature map output by the global adapted ViT backbone, which can also be regarded as coarse-grained patch-level features. However, in order to achieve promising performance improvements through (local matching) re-ranking, more fine-grained dense local features are required. To achieve this, we propose the local adaptation, which is achieved by an up-sampling module after the ViT backbone, as shown in Fig. 3. To be specific, this module consists of two up-convolutional (up-conv) layers and a ReLU layer in the middle. The height and width of the feature map will approximately double after passing through each up-conv layer, while the channel dimension will be reduced. Finally, for the input image of size 224$\times$224 pixels, this module adjusts the 16$\times$16$\times$1024-dim feature map output by ViT-L/14 backbone to 61$\times$61$\times$128-dim, and performs L2 normalization in the channel dimension (intraL2) to yield dense local features, i.e., a dense 61$\times$61 grid of 128-dim local features $\bm{f}^{l}$. Formally, it can be represented as
@@ 3.4 Local Matching for Re-ranking
Our two-stage place retrieval pipeline is shown in Fig. 3. After obtaining the global features and local features, we first compute L2 distance to perform the similarity search in the global feature space over the database to get the top-k most similar candidate images. For the local features matching between the query image $q$ and a candidate image $c$, we search for mutual nearest neighbor matches by cross-matching. Since local features are L2 normalized, the inner product equivalent to cosine similarity is used to measure local feature similarity. That is
$\bm{s}_{qc}(i,j)=\bm{f}^{l}_{q}(i)\cdot\bm{f}^{l}_{c}(j)\quad i,j\in\{1,2,...,N^{\prime}\}\quad(N^{\prime}=61\times 61),$ (7)
@@
That is, the $u$-th feature in image $q$ and the $v$-th feature in image $c$ are the best matches for each other. The obtained nearest neighbor matches set in previous VPR work (Hausler et al., 2021; Wang et al., 2022a) exists a large number of false matches. So they apply geometric verification (e.g. RANSAC (Fischler & Bolles, 1981)) to remove outliers (i.e. false matches) and use the number of inliers as the image similarity score, which is time-consuming. Due to the powerful representation ability of the foundation model and the superiority of the ViT in capturing long-distance feature dependencies, the number of false matches in this work is actually not enough to affect the performance of re-ranking. So we directly use the number of matches (i.e., $|\mathcal{M}|$) as the image similarity score for re-ranking.
@@ 4.1 Datasets and Performance Evaluation
Several VPR benchmark datasets mainly including Tokyo24/7, MSLS, and Pitts30k are used in our experiments. Table 1 summarizes their main information. Tokyo24/7 (Torii et al., 2015) includes about 76k database images and 315 query images captured from urban scenes with drastic illumination changes. Mapillary Street-Level Sequences (MSLS) (Warburg et al., 2020) consists of more than 1.6 million images collected in urban, suburban and natural scenes over 7 years. We assess models on both MSLS-val and MSLS-challenge (an online test set without released labels) sets. Pittsburgh (Pitts30k) (Torii et al., 2013) contains 30k reference images and 24k query images in the train, val and test sets, and exists severe viewpoint changes. More details are in Appendix J.
@@ 4.2 Implementation Details
We use the DINOv2 based on ViT-L/14 as the foundation model and conduct all experiments on an NVIDIA GeForce RTX 3090 GPU using PyTorch. Fed a 224×224 image, the model produces a 1024-dim global feature and a dense grid of 128-dim local features. The bottleneck ratio of the adapters in ViT blocks is 0.5 and the scaling factor $s$ in Eq. 4 is set to 0.2. We use 3×3 up-conv with stride=2 and padding=1 in the local adaptation module. The output channels of the first and second up-conv layers are 256 and 128, respectively. Following other two-stage methods, we re-rank the top-100 candidates to yield final results. We train our models using the Adam optimizer with the learning rate set as 0.00001 and batch size set as 4. When the R@5 on the validation set does not have improvement within 3 epochs, the training is terminated. For MSLS, we set an epoch as passing 30k queries, whereas Pitts30k is passing 5k queries. In model training, we define the potential positive images as the reference images that are within 10 meters from the query image, while the definite negative images are those further than 25 meters. Two hard negative images from 1000 randomly chosen definite negatives are used in the triplet loss. We empirically set the margin $m=0.1$ in Eq. 9, the weight $\lambda=1$ in Eq. 11. For the experiments in Section 4.3 (i.e. comparisons with other methods), we fine-tune our model on MSLS-train to test on MSLS-val/challenge, and further fine-tune it on Pitts30k-train to test on Pitts30k-test and Tokyo24/7 (as in $R^{2}$ormer (Zhu et al., 2023)). For the ablation experiments in other sections, the models tested on Pitts30k-test and Tokyo24/7 are directly fine-tuned on Pitts30k-train by default.
@@ 4.3 Comparisons with State-of-the-Art Methods
In this section, we compare the proposed SelaVPR method with several state-of-the-art (SOTA) VPR methods, including four one-stage methods using global feature retrieval: NetVLAD (Arandjelovic et al., 2016), SFRS (Ge et al., 2020), CosPlace (Berton et al., 2022a) and MixVPR (Ali-Bey et al., 2023), as well as five two-stage methods with re-ranking: SP-SuperGlue (DeTone et al., 2018; Sarlin et al., 2020), Patch-NetVLAD (Hausler et al., 2021), TransVPR (Wang et al., 2022a), StructVPR (Shen et al., 2023) and $R^{2}$ormer (Zhu et al., 2023). The details of these methods are in Appendix K. Note that CosPlace and MixVPR are trained on individually constructed large-scale datasets. TransVPR and $R^{2}$ormer are transformer-based methods. $R^{2}$ormer ranked first in the leaderboard of the MSLS place recognition challenge before our method was proposed. We also show the global retrieval result without re-ranking using the proposed SelaVPR, and denote it as SelaVPR(global). The quantitative results are shown in Table 2. Since the code for StructVPR has not yet been released, we use the results reported in the paper (Shen et al., 2023). The proposed method achieves the best R@1/R@5/R@10 on all datasets.
When using only global features for direct retrieval, SelaVPR(global) significantly outperforms other one-stage methods on all datasets for R@5 and R@10, including SFRS using complex training strategies and CosPlace/MixVPR trained on purpose-built large-scale datasets. SelaVPR(global) also outperforms all other two-stage methods on Tokyo24/7, MSLS-val, and MSLS-challenge for R@5 and R@10. This fully demonstrates that adapting the foundation model can provide powerful feature representation, which is a novel way to achieve SOTA one-stage VPR. Although SelaVPR(global) does not achieve very good R@1 on Tokyo24/7 where the lighting changes drastically, the complete SelaVPR method outperforms other methods by a large margin after local feature re-ranking. This illustrates the necessity of using local matching (local adaptation) for VPR in extreme environments. The complete SelaVPR method significantly outperforms SOTA methods on Tokyo24/7 and Pitts30k with absolute R@1 improvement of 5.4% and 1.3% respectively. Meanwhile, it also ranks 1st on the leaderboard of MSLS place recognition challenge (see Appendix L). Comparisons to SOTA methods on more datasets are in Appendix F (SelaVPR achieves near-perfect results on St. Lucia). Fig. 4 qualitatively demonstrates that our approach is highly robust in challenging scenes. Benefiting from the visual foundation model and sensible adaptation, SelaVPR achieves absolute performance advantages on a variety of datasets without complex training strategies or purpose-built large-scale training datasets.
@@
Efficiency is another metric for evaluating the VPR methods. In Table 3, we compare the runtime (feature extraction time and matching/retrieval time) of our method with other two-stage methods on Pitts30k-test. Patch-NetVLAD-p and TransVPR are representative two-stage methods using RANSAC for spatial verification in re-ranking. SP-SuperGlue uses neural networks to match local features. TransVPR is fast at extracting features, while SelaVPR is slower (but faster than other methods) due to the use of the ViT/L backbone. However, for matching/retrieval, since our method does not require time-consuming spatial verification in re-ranking, its runtime is less than 3% of TransVPR and only about 1% of SP-SuperGlue and Patch-NetVLAD-p. The total runtime of our method is less than 4% of TransVPR. Although Patch-NetVLAD-s uses a Rapid Spatial Scoring method for fast verification, it is still significantly slower than ours. Fig. 5 simultaneously shows the total runtime and R@1 on Pitts30k. Due to the absolute advantages in both performance and efficiency, our SelaVPR method is able to pave the way for real-world large-scale VPR applications.